How do I?
Provide an example that uses the substitution method to solve two linear equations.
I do not understand at all

Answer by Alan3354(69443)

x + y = 10
2x - y = 4 --> y = 2x - 4
-------------
Sub for y in the 1st eqn
x + (2x+4) = 10
Solve for x